# A Survey of Reinforcement Learning For Economics

### 09 Mar 2026 v1

@@ 4.3.2 Why Off-Policy Learning Diverges
In the off-policy setting, samples come from a behavior distribution $\mu\neq d^{\pi}$. The projection now minimizes error under $\mu$, but the Bellman operator $T^{\pi}$ still contracts in the $d^{\pi}$-norm. The two operators measure distance in different norms. The projection is no longer orthogonal in the $d^{\pi}$-norm; it is oblique. Unlike orthogonal projections, oblique projections can expand distances: $\|\Pi_{\mu}\|_{d^{\pi}}$ can exceed 1. If $\|\Pi_{\mu}\|_{d^{\pi}}>1/\gamma$, the expansion from projection overwhelms the $\gamma$-contraction from the Bellman operator, and the fixed-point iteration diverges.
@@
This divergence is not overfitting. Overfitting occurs when the approximator memorizes training data at the expense of generalization; collecting more data helps. Divergence means the parameter vector $\theta$ grows without bound, producing arbitrarily large value estimates that bear no relation to the true values. More data does not help; the algorithm itself is unstable. The distinction matters because the remedies are entirely different: regularization and early stopping address overfitting, while the deadly triad requires structural changes to the algorithm.
@@
Baird (1995) constructed a six-state star MDP that makes this failure concrete. All rewards are zero, so the true value is $V^{*}(s)=0$ for every state. A lookup table learns this immediately. The MDP has a star topology: states 1 through 5 each transition to state 6, and state 6 transitions to itself. Linear function approximation uses a shared weight $w_{1}$ across all states plus a state-specific weight, with $V(s)=2w_{1}+w_{s}$ for $s\in\{1,\ldots,5\}$ and $V(6)=2w_{1}-w_{6}$. Training samples all transitions equally often (uniform distribution, not $d^{\pi}$). The dynamics are as follows. [^56] Baird (1995) also presents an MDP variant with two actions per state, demonstrating that Q-learning diverges under the same mechanism. The key insight is identical: shared parameters create cross-state coupling that uniform sampling cannot counterbalance. When $V(6)$ is large and positive, the TD target $\gamma V(6)$ exceeds $V(s)$ for states 1 through 5, producing positive TD errors that push $w_{1}$ upward. At state 6, the TD target is $\gamma V(6)<V(6)$, producing a negative TD error that pushes $w_{1}$ downward. But states 1 through 5 are each visited as often as state 6, so $w_{1}$ receives five upward pushes for every one downward push. The shared weight diverges to $\infty$. The on-policy distribution would concentrate mass on state 6 (the absorbing state), counterbalancing the upward pressure; uniform sampling destroys this balance. [^57] The gradient of the mean-squared Bellman error $\|Q-TQ\|^{2}$ would correct for this imbalance, but computing an unbiased gradient requires two independent next-state samples from the same $(s,a)$, since $\nabla\mathbb{E}[(Q-\mathbb{E}[r+\gamma V(s^{\prime})])^{2}]$ involves the product $\mathbb{E}[\cdot]\cdot\nabla\mathbb{E}[\cdot]$. A single sample yields a biased estimate ($\mathbb{E}[XY]\neq\mathbb{E}[X]\mathbb{E}[Y]$). This “double-sampling” requirement is impractical in most environments (Baird, 1995). Practitioners instead use semi-gradient TD, which treats the bootstrap target as a constant and finds the projected fixed point $\Pi T^{\pi}\theta^{*}$ rather than the true Bellman fixed point. It is this semi-gradient structure that makes off-policy TD inherently vulnerable to the projection mismatch.
@@ 4.4.1 The Policy Gradient Theorem
The policy gradient theorem sidesteps this entirely via a likelihood ratio (or score function) trick. [^60] The score function $\nabla_{\theta}\log\pi_{\theta}(a|s)$ is the same mathematical object that appears in the Cramér-Rao bound and the score test in maximum likelihood estimation. The policy gradient is a covariance: $\nabla J=\text{Cov}_{d^{\pi_{\theta}}\times\pi_{\theta}}(\nabla\log\pi_{\theta},Q^{\pi_{\theta}})$, measuring how sensitive the log-likelihood of the policy is to parameter changes, weighted by action quality. The theorem transforms a sensitivity analysis problem (how does the system evolve?) into a simpler expectation problem (what is the correlation between the score $\nabla\log\pi$ and the value $Q$?). The gradient can be written as an expectation under the current policy, weighted by action-values, without requiring $\nabla_{\theta}d^{\pi_{\theta}}$. The transition dynamics $P(s^{\prime}|s,a)$ do not appear; the gradient is estimable via sample averages from trajectories alone.
@@ 4.4.3 Natural Policy Gradient and Gradient Domination
Standard gradient descent treats all parameter directions equally. But small changes in $\theta$ can cause large changes in the policy distribution $\pi_{\theta}$. The natural policy gradient (Kakade, 2001), building on the natural gradient framework of Amari (1998), accounts for this curvature by preconditioning with the Fisher information matrix. [^62] The Fisher information $F(\theta)=\mathbb{E}[\nabla\log p\nabla\log p^{\top}]$ measures curvature of the log-likelihood and appears in the Cramér-Rao bound. Here it measures how policy distributions change with parameters. The relationship between NPG and standard PG parallels that between GLS and OLS in econometrics: both account for the covariance structure of the estimating equations, yielding efficiency gains when the covariance is non-spherical.
@@
The RL objective $J(\theta)$ is non-convex in $\theta$. For economists trained to distrust gradient methods on non-convex objectives, the natural concern is convergence to spurious local optima. For softmax policies, this concern is unfounded. The landscape is “benign” in a precise sense. Agarwal et al. (2021) prove that $J(\theta)$ satisfies a gradient domination condition (also called Polyak-Łojasiewicz, or PL). The PL condition has the same functional form as the strong convexity condition for guaranteeing linear convergence of gradient descent, but it applies to non-convex functions: whenever $\|\nabla J(\theta)\|$ is small, the policy must be near-optimal. Formally, the sub-optimality $J(\pi^{*})-J(\pi_{\theta})$ is bounded by a constant times $\|\nabla J(\theta)\|^{2}$. The implication is immediate: any point where the gradient vanishes is globally optimal. The non-convex landscape has no false peaks, no spurious local maxima. Gradient ascent cannot get trapped.
@@
Mei et al. (2020) sharpen this result for softmax parameterization, proving explicit convergence rates. [^63] However, “no spurious local optima” does not imply “easy optimization.” The landscape is dominated by vast plateaus (saddle points) where gradients vanish. Without sufficient exploration, the probability of visiting relevant states decays exponentially with the horizon, rendering the gradient exponentially small. Global convergence requires the starting distribution to have adequate coverage relative to the optimal policy’s visitation distribution, formalized as the “distribution mismatch coefficient” by Agarwal et al. (2021). The Natural Policy Gradient addresses this by preconditioning with the Fisher Information Matrix, making the update direction covariant: invariant to invertible linear transformations of the parameter space. This standardizes units across parameters, preventing stalling on plateaus caused by poor parameter scaling. Li et al. (2022) make this quantitatively precise: vanilla softmax policy gradient requires iterations doubly exponential in the effective horizon $1/(1-\gamma)$ because score functions are exponentially small in directions corresponding to suboptimal actions.
@@
The Natural Policy Gradient achieves more: dimension-free convergence. Standard gradient ascent on $J(\theta)$ has a convergence rate that depends on the smoothness constant, which scales with $|\mathcal{S}|$. NPG circumvents this by preconditioning with the Fisher information matrix $F(\theta)^{-1}$. The mechanism is that the state-visitation distribution $d^{\pi_{\theta}}(s)$ appears in both $\nabla J(\theta)$ and $F(\theta)$; when computing $F^{-1}\nabla J$, these terms cancel analytically. The resulting update rule is equivalent to soft policy iteration and converges at rate $O(1/(1-\gamma)^{2}\epsilon)$, independent of $|\mathcal{S}|$ and $|\mathcal{A}|$ (Xiao, 2022).
@@ 4.4.4 Trust Region Methods
NPG requires computing and inverting the Fisher information matrix $F(\theta)$, which scales as $O(d^{2})$ in parameters and is impractical for neural networks. TRPO (Schulman et al., 2015) approximates the natural gradient using conjugate gradient methods [^64] Conjugate gradient is an iterative method for solving linear systems $Ax=b$ without forming $A$ explicitly, requiring only matrix-vector products $Av$. With $k$ iterations it costs $O(kd)$ versus $O(d^{3})$ for direct inversion, making it feasible for neural networks with millions of parameters. without forming $F$ explicitly, and enforces trust regions via line search. [^65] Trust region methods build on the performance difference lemma: for any two policies $\pi$ and $\pi^{\prime}$, $J(\pi^{\prime})-J(\pi)=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi^{\prime}}}[\sum_{a}\pi^{\prime}(a|s)A^{\pi}(s,a)]$, where $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$ is the advantage function. This identity bounds how much policy improvement is possible and motivates constraining updates to regions where advantage estimates remain accurate (Kakade, 2002). Shani et al. (2020) prove convergence for adaptive trust region methods that adjust the constraint radius dynamically. PPO (Schulman et al., 2017) simplifies further by replacing the hard KL constraint with a clipped surrogate objective, trading theoretical guarantees for computational simplicity.

### 15 Mar 2026 v2

@@ 5.3.2 Why Off-Policy Learning Diverges
In the off-policy setting, samples come from a behavior distribution $\mu\neq d^{\pi}$. The projection now minimizes error under $\mu$, but the Bellman operator $T^{\pi}$ still contracts in the $d^{\pi}$-norm. The two operators measure distance in different norms. The projection is no longer orthogonal in the $d^{\pi}$-norm; it is oblique. Unlike orthogonal projections, oblique projections can expand distances, with $\|\Pi_{\mu}\|_{d^{\pi}}$ exceeding 1 in the worst case. If $\|\Pi_{\mu}\|_{d^{\pi}}>1/\gamma$, the expansion from projection overwhelms the $\gamma$-contraction from the Bellman operator, and the fixed-point iteration diverges.
@@
This divergence is not overfitting. Overfitting occurs when the approximator memorizes training data at the expense of generalization; collecting more data helps. Divergence means the parameter vector $\theta$ grows without bound, producing arbitrarily large value estimates that bear no relation to the true values. More data does not help; the algorithm itself is unstable. The distinction matters because the remedies are entirely different. Regularization and early stopping address overfitting, while the deadly triad requires structural changes to the algorithm.
@@
Baird (1995) constructed a six-state star MDP that makes this failure concrete. All rewards are zero, so the true value is $V^{*}(s)=0$ for every state. A lookup table learns this immediately. The MDP has a star topology: states 1 through 5 each transition to state 6, and state 6 transitions to itself. Linear function approximation uses a shared weight $w_{1}$ across all states plus a state-specific weight, with $V(s)=2w_{1}+w_{s}$ for $s\in\{1,\ldots,5\}$ and $V(6)=2w_{1}-w_{6}$. Training samples all transitions equally often (uniform distribution, not $d^{\pi}$). The dynamics are as follows. [^68] Baird (1995) also presents an MDP variant with two actions per state, demonstrating that Q-learning diverges under the same mechanism. The key insight is identical: shared parameters create cross-state coupling that uniform sampling cannot counterbalance. When $V(6)$ is large and positive, the TD target $\gamma V(6)$ exceeds $V(s)$ for states 1 through 5, producing positive TD errors that push $w_{1}$ upward. At state 6, the TD target is $\gamma V(6)<V(6)$, producing a negative TD error that pushes $w_{1}$ downward. But states 1 through 5 are each visited as often as state 6, so $w_{1}$ receives five upward pushes for every one downward push. The shared weight diverges to $\infty$. The on-policy distribution would concentrate mass on state 6 (the absorbing state), counterbalancing the upward pressure; uniform sampling destroys this balance. [^69] The gradient of $\|Q-TQ\|^{2}$ requires two independent next-state samples from the same $(s,a)$, since $\nabla\mathbb{E}[(Q-\mathbb{E}[r+\gamma V(s^{\prime})])^{2}]$ involves $\mathbb{E}[\cdot]\cdot\nabla\mathbb{E}[\cdot]$ and $\mathbb{E}[XY]\neq\mathbb{E}[X]\mathbb{E}[Y]$. This “double-sampling” requirement is impractical (Baird, 1995), so practitioners use semi-gradient TD, treating the bootstrap target as a constant. This semi-gradient structure makes off-policy TD vulnerable to projection mismatch.
@@
Figure 7: Geometry of the projected Bellman operator in $\mathbb{R}^{2}$. The gray line is the function approximation subspace span($\Phi$); the blue arrow is $TV$, the Bellman update. (a) On-policy: the orthogonal projection $\Pi_{\mu}$ drops TVTV perpendicularly onto the subspace, preserving the contraction. (b) Off-policy: the oblique projection $\Pi_{\nu}$ (under the behavior distribution) reaches a point further from the origin than TVTV itself, causing expansion.
@@ 5.4.1 The Policy Gradient Theorem
The policy gradient theorem sidesteps this entirely via a likelihood ratio (or score function) trick. [^72] The score function $\nabla_{\theta}\log\pi_{\theta}(a|s)$ is the same mathematical object that appears in the Cramér-Rao bound and the score test in maximum likelihood estimation. The policy gradient is a covariance, $\nabla J=\text{Cov}_{d^{\pi_{\theta}}\times\pi_{\theta}}(\nabla\log\pi_{\theta},Q^{\pi_{\theta}})$, measuring how sensitive the log-likelihood of the policy is to parameter changes, weighted by action quality. The theorem transforms a sensitivity analysis problem (how does the system evolve?) into a simpler expectation problem (what is the correlation between the score $\nabla\log\pi$ and the value $Q$?). The gradient can be written as an expectation under the current policy, weighted by action-values, without requiring $\nabla_{\theta}d^{\pi_{\theta}}$. The transition dynamics $P(s^{\prime}|s,a)$ do not appear; the gradient is estimable via sample averages from trajectories alone.
@@ 5.4.3 Natural Policy Gradient and Gradient Domination
Standard gradient descent treats all parameter directions equally. But small changes in $\theta$ can cause large changes in the policy distribution $\pi_{\theta}$. The natural policy gradient (Kakade, 2001), building on the natural gradient framework of Amari (1998), accounts for this curvature by preconditioning with the Fisher information matrix. [^74] The Fisher information $F(\theta)=\mathbb{E}[\nabla\log p\nabla\log p^{\top}]$ measures curvature of the log-likelihood and appears in the Cramér-Rao bound. Here it measures how policy distributions change with parameters. The relationship between NPG and standard PG parallels that between Fisher scoring and gradient ascent in MLE. Both precondition with the inverse Fisher information matrix $F(\theta)^{-1}$, achieving parameterization invariance and quadratic convergence near the optimum.
@@
The RL objective $J(\theta)$ is non-convex in $\theta$. For economists trained to distrust gradient methods on non-convex objectives, the natural concern is convergence to spurious local optima. For tabular softmax policies (one free parameter per state-action pair), this concern is unfounded. The landscape is “benign” in a precise sense. Agarwal et al. (2021a) prove that $J(\theta)$ satisfies a gradient domination condition (also called Polyak-Łojasiewicz, or PL). The PL condition has the same functional form as the strong convexity condition for guaranteeing linear convergence of gradient descent, but it applies to non-convex functions: whenever $\|\nabla J(\theta)\|$ is small, the policy must be near-optimal. Formally, the sub-optimality $J(\pi^{*})-J(\pi_{\theta})$ is bounded by a constant times $\|\nabla J(\theta)\|^{2}$. The implication is immediate: any point where the gradient vanishes is globally optimal. The non-convex landscape has no false peaks, no spurious local maxima. Gradient ascent cannot get trapped.
@@
Mei et al. (2020) sharpen this result for softmax parameterization, proving explicit convergence rates. These guarantees are specific to the tabular parameterization. With function approximation, the PL condition does not hold. Agarwal et al. (2021a, Theorem 6.2) show that NPG with log-linear or smooth policy classes (including neural networks) converges to a neighborhood of the optimum whose radius depends on the approximation error of the policy class, not to the global optimum itself. [^75] However, “no spurious local optima” does not imply “easy optimization.” The landscape is dominated by vast plateaus (saddle points) where gradients vanish. Without sufficient exploration, the probability of visiting relevant states decays exponentially with the horizon, rendering the gradient exponentially small. Global convergence requires the starting distribution to have adequate coverage relative to the optimal policy’s visitation distribution, formalized as the “distribution mismatch coefficient” by Agarwal et al. (2021a). The Natural Policy Gradient addresses this by preconditioning with the Fisher Information Matrix, making the update direction covariant: invariant to invertible linear transformations of the parameter space. This standardizes units across parameters, preventing stalling on plateaus caused by poor parameter scaling. Li et al. (2022) make this quantitatively precise: vanilla softmax policy gradient requires iterations doubly exponential in the effective horizon $1/(1-\gamma)$ because score functions are exponentially small in directions corresponding to suboptimal actions.
@@
In the tabular setting, NPG achieves more: dimension-free convergence. Standard gradient ascent on $J(\theta)$ has a convergence rate that depends on the smoothness constant, which scales with $|\mathcal{S}|$. NPG circumvents this by preconditioning with the Fisher information matrix $F(\theta)^{-1}$. The mechanism is that the state-visitation distribution $d^{\pi_{\theta}}(s)$ appears in both $\nabla J(\theta)$ and $F(\theta)$; when computing $F^{-1}\nabla J$, these terms cancel analytically. The resulting update rule is equivalent to soft policy iteration and converges at rate $O(1/(1-\gamma)^{2}\epsilon)$, independent of $|\mathcal{S}|$ and $|\mathcal{A}|$ (Xiao, 2022).
@@ 5.4.4 Trust Region Methods
NPG requires computing and inverting the Fisher information matrix $F(\theta)$, which scales as $O(d^{2})$ in parameters and is impractical for neural networks. TRPO (Schulman et al., 2015) approximates the natural gradient using conjugate gradient methods [^76] Conjugate gradient is an iterative method for solving linear systems $Ax=b$ without forming $A$ explicitly, requiring only matrix-vector products $Av$. With $k$ iterations it costs $O(kd)$ versus $O(d^{3})$ for direct inversion, making it feasible for neural networks with millions of parameters. without forming $F$ explicitly, and enforces trust regions via line search. [^77] Trust region methods build on the performance difference lemma: for any two policies $\pi$ and $\pi^{\prime}$, $J(\pi^{\prime})-J(\pi)=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi^{\prime}}}[\sum_{a}\pi^{\prime}(a|s)A^{\pi}(s,a)]$, where $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$ is the advantage function. This identity bounds how much policy improvement is possible and motivates constraining updates to regions where advantage estimates remain accurate (Kakade, 2002). Shani et al. (2020) prove convergence for adaptive trust region methods that adjust the constraint radius dynamically. PPO (Schulman et al., 2017) simplifies further by replacing the hard KL constraint with a clipped surrogate objective, trading theoretical guarantees for computational simplicity. The geometric foundation of trust region methods lies in information geometry. The space of policies $\{\pi_{\theta}:\theta\in\mathbb{R}^{d}\}$ forms a statistical manifold, and the natural distance between two nearby policies is the KL divergence, not the Euclidean distance between their parameters (Amari, 1998). To second order, $\mathrm{KL}(\pi_{\theta}\|\pi_{\theta+\Delta\theta})\approx\frac{1}{2}\Delta\theta^{\top}F(\theta)\Delta\theta$, where $F(\theta)$ is the Fisher information matrix. Two parameter vectors $\theta$ and $\theta^{\prime}$ that are far apart in Euclidean distance may correspond to nearly identical distributions, while nearby parameters may produce radically different policies. The natural gradient corrects for this by measuring steepest ascent in KL-divergence rather than Euclidean norm. Figure 8 illustrates the distinction: on the policy manifold, the Euclidean gradient $\nabla_{\theta}J$ points in a direction that ignores curvature, while the natural gradient $F^{-1}\nabla_{\theta}J$ follows the manifold’s intrinsic geometry toward the optimum.
Figure 8: Information geometry of the natural policy gradient. Left: the policy manifold $\mathcal{M}$ with Euclidean gradient $\nabla_{\theta}J$ (red) and natural gradient $F^{-1}\nabla_{\theta}J$ (green) from the current iterate $\pi_{\theta_{\mathrm{old}}}$ toward the optimal policy $\pi^{*}$. Right: tangent plane at $\theta_{\mathrm{old}}$ showing the Euclidean unit ball $\|\Delta\theta\|_{2}\leq\varepsilon$ (red) and the KL unit ball $\Delta\theta^{\top}F\Delta\theta\leq\delta$ (green), with the respective steepest-ascent directions.
TRPO formalizes this insight as a constrained optimization problem. At each iteration, TRPO maximizes the importance-weighted surrogate
$\max_{\theta}\;L(\theta)=\mathbb{E}_{s\sim d^{\pi_{\theta_{\mathrm{old}}}}}\left[\sum_{a}\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{\mathrm{old}}}(a|s)}A^{\pi_{\theta_{\mathrm{old}}}}(s,a)\right]\quad\text{s.t.}\quad\mathrm{KL}(\pi_{\theta_{\mathrm{old}}}\|\pi_{\theta})\leq\delta,$ (50)
where $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$ is the advantage function. Linearizing $L(\theta)$ around $\theta_{\mathrm{old}}$ and applying the quadratic KL approximation yields a Lagrangian whose closed-form solution is
$\theta_{\mathrm{new}}=\theta_{\mathrm{old}}+\sqrt{\frac{2\delta}{g^{\top}F^{-1}g}}\,F^{-1}g,\qquad g=\nabla_{\theta}L(\theta)\big|_{\theta_{\mathrm{old}}}.$ (51)
This is precisely the natural gradient direction, scaled so that the step saturates the KL budget $\delta$. The step size is determined entirely by the trust region geometry, not by a learning rate hyperparameter. In practice, TRPO solves the linear system $Fv=g$ via conjugate gradient and performs a backtracking line search to enforce the KL constraint exactly. [^78] Conjugate gradient solves $Fv=g$ iteratively using only matrix-vector products $Fv$, which can be computed via automatic differentiation without forming $F$ explicitly. With $k$ iterations it costs $O(kd)$ versus $O(d^{3})$ for direct inversion, making it feasible for neural networks with millions of parameters.
The theoretical guarantee underlying TRPO is a majorization-minimization (MM) argument. The surrogate $L(\theta)$ is a local lower bound on $J(\theta)$ that is tight at $\theta_{\mathrm{old}}$: $L(\theta_{\mathrm{old}})=J(\theta_{\mathrm{old}})$ and $L(\theta)\leq J(\theta)$ within the trust region. [^79] The bound follows from the performance difference lemma: $J(\pi^{\prime})-J(\pi)=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi^{\prime}}}[\sum_{a}\pi^{\prime}(a|s)A^{\pi}(s,a)]$. Replacing $d^{\pi^{\prime}}$ with $d^{\pi}$ introduces error controlled by the KL divergence between the two policies (Kakade, 2002). Maximizing $L$ within the trust region therefore guarantees monotonic improvement: $J(\theta_{\mathrm{new}})\geq L(\theta_{\mathrm{new}})\geq L(\theta_{\mathrm{old}})=J(\theta_{\mathrm{old}})$. This is the same pattern as the EM algorithm in statistics, where the E-step constructs a surrogate (the ELBO) and the M-step maximizes it. [^80] In EM, $Q(\theta|\theta^{(t)})$ lower-bounds the log-likelihood and is tight at $\theta^{(t)}$. Each M-step guarantees $\ell(\theta^{(t+1)})\geq Q(\theta^{(t+1)}|\theta^{(t)})\geq Q(\theta^{(t)}|\theta^{(t)})=\ell(\theta^{(t)})$. The TRPO bound has the same structure with $L$ playing the role of $Q$ and $J$ playing the role of $\ell$. Shani et al. (2020) prove convergence for adaptive trust region methods that adjust $\delta$ dynamically. Figure 9 illustrates this mechanism: each surrogate is a lower bound that touches $J$ at the current iterate, and sequential maximization produces monotonically improving iterates converging to $\theta^{*}$.
Figure 9: Majorization-minimization interpretation of trust region updates. Left: the surrogate $L(\theta|\theta_{\mathrm{old}})$ (dashed) lower-bounds $J(\theta)$ (solid) and is tight at $\theta_{\mathrm{old}}$; the trust region (shaded) constrains the step. The gap between $L(\theta_{\mathrm{new}})$ and $J(\theta_{\mathrm{new}})$ is the guaranteed improvement. Right: iterative $M$ convergence from $\theta_{0}$ through four surrogates (dashed, colored by iteration) to $\theta^{*}$.
PPO (Schulman et al., 2017) replaces the hard KL constraint with a clipped surrogate objective. Let $r_{t}(\theta)=\pi_{\theta}(a_{t}|s_{t})/\pi_{\theta_{\mathrm{old}}}(a_{t}|s_{t})$ denote the importance ratio. PPO maximizes
$L^{\mathrm{clip}}(\theta)=\mathbb{E}_{t}\left[\min\!\big(r_{t}(\theta)A_{t},\;\mathrm{clip}(r_{t}(\theta),1-\varepsilon,1+\varepsilon)A_{t}\big)\right],$ (52)
where $\varepsilon$ (typically 0.1–0.2) bounds the ratio. When $A_{t}>0$, clipping prevents $r_{t}$ from exceeding $1+\varepsilon$; when $A_{t}<0$, it prevents $r_{t}$ from falling below $1-\varepsilon$. The resulting feasible region is not an ellipsoid in parameter space but rather a non-convex set determined by the ratio constraint at each sampled state-action pair. PPO requires no Fisher information computation and uses only first-order gradients, making it the dominant method in large-scale applications including RLHF (Section 12). Figure 10 illustrates all three mechanisms in the LQC monetary policy setting, where the non-ellipsoidal PPO feasible region is visible in contrast to TRPO’s KL ellipse.
The trust region framework connects naturally to econometric optimization. The Levenberg-Marquardt algorithm for nonlinear least squares uses a similar trust region mechanism, interpolating between gradient descent and Gauss-Newton steps. [^81] Levenberg-Marquardt solves $\min_{\theta}\|r(\theta)\|^{2}$ by adding a damping term $\lambda I$ to the Gauss-Newton Hessian approximation $J^{\top}J$, which is equivalent to constraining the step to a trust region whose radius decreases with $\lambda$. TRPO replaces $J^{\top}J$ with the Fisher information matrix $F(\theta)$. More broadly, the Fisher information matrix that defines TRPO’s trust region is the same object that appears in the Cramér-Rao bound: it measures the statistical precision of the policy parameterization. The natural gradient adapts step sizes to this precision, taking large steps in well-identified directions and small steps where the data provide little information about the policy.
Figure 10: Trust region methods in the LQC monetary policy setting. A central bank learns a Taylor rule $u_{t}=-(\theta_{1}x_{1t}+\theta_{2}x_{2t})$ mapping output gap $x_{1}$ and inflation gap $x_{2}$ to an interest rate instrument. Left: policy contour lines in state space for the current iterate $\theta_{\mathrm{old}}$, optimal weights $\theta^{*}$, and the unconstrained gradient step $\theta_{\mathrm{bad}}$, with phase arrows showing closed-loop dynamics under $\theta_{\mathrm{old}}$. Center: expected return $J(\theta_{1},\theta_{2})$ in parameter space; hatching marks the unstable region. The KL trust region ellipse bounds the TRPO step; the unconstrained gradient step overshoots into the unstable region. Right: TRPO feasible region (KL ellipse) and PPO feasible region (50% ratio-clip band over 200 sampled state-action pairs) overlaid on $J(\theta)$, with the respective constrained steps marked.